OpenStudy (anonymous):

find the value of:3(8+3)-4

OpenStudy (anonymous):

Distribute the 3 into the parenthesis, then add it all together.

OpenStudy (anonymous):

3(8+3)-4 = (8 x 3 ) + (3 x 3) - 4 = 24 + 9 - 4

OpenStudy (anonymous):

= 29

OpenStudy (anonymous):

is not 30

OpenStudy (anonymous):

add 8 and 3, then multiple that by 3 then subtract 4 = 29

OpenStudy (anonymous):

i said 29

OpenStudy (anonymous):

no

OpenStudy (anonymous):

thanks mustafa102

OpenStudy (anonymous):

No problem